OpenStudy (anonymous):

Which line is the same as [x, y]=[1, -8] + t[ 4,-3]? A [x, y] =[2, 4] +t[-4, 3] B [x, y] =[13, 17]+t[8, -6]

OpenStudy (anonymous):

And how do you figure it out?

OpenStudy (anonymous):

you could just determine the line equation y(x), through substitution

OpenStudy (anonymous):

@completeidiot how ?

OpenStudy (anonymous):

[x, y]=[1, -8] + t[ 4,-3] x=1+4t y=-8-3t solve for t for one of the equations, then substitute

OpenStudy (anonymous):

by elimination ? @completeidiot

OpenStudy (anonymous):

look for the coefficients of parallel vector. if they are same means either same line or parallel

OpenStudy (anonymous):

@sami-21 what do you mean ?

OpenStudy (anonymous):

look the coefficients of parameter t . each line in space has parallel vector to describe it . so if the vectors are same then lts same .

OpenStudy (anonymous):

in other words i am saying compare following x=1+4t y=-8-3t with A x=2-4t y=4+3t and B x=13+8t y=17-6t

OpenStudy (anonymous):

did you get it ?

OpenStudy (anonymous):

im trying to understand it :S

OpenStudy (anonymous):

well, let me show you how B is same as the given line. each line in the space requires one point and vector to completely define it the coefficients of the t represents that vector so if the coefficients of the t are same or scalar multiple of some number then lines are definitely same . so the vector parallel to given line is (coefficients of t ) v1=<4,-3> vector paralel to line A (coefficients of t ) vA=<-4,3> vector parallel to line B (coefficients of t ) vB=<8,-6> you can see that the vetor Vb parallel to line B is 2 times the vector parallel to given line vB=2v1 it means given line and line |B are identical. Hope its clear now .

OpenStudy (anonymous):

Like how do i compare the parametric equations of A and B. Is there something i sub in ?

OpenStudy (anonymous):

no i meant solve for t... x=1+4t x-1=4t t=(x-1)/4 y=-8-3t y=-8-3((x-1)/4)